### v1

@@ Appendix D Omitted Proofs of Section 6
Refer to Figure 6 for an illustration of our construction. We first show inductively that the edges on the left side are assigned to the same priority queues as their (adjacent) synchronization edges, and then we show the same for the edges on the right side.
Let $a_{1}$ be the edge with the leftmost left endpoint corresponding to a vertex $a\in S$. The $k-1$ heavy edges below $a_{1}$ require $k-1$ priority queues, none of which can be used by $a_{1}$ or its synchronization edge since both form a forbidden pattern with each of the heavy edges. Therefore, $a_{1}$ and its synchronization edge are assigned to the only remaining priority queue.
Let $b_{1}$ be the edge with the second leftmost left endpoint corresponding to a vertex $b\in S$. The $k-2$ heavy edges below $b_{1}$ require $k-2$ priority queues, none of which can be used by $b_{1}$ or its synchronization edge. Moreover, $a_{1}$ and its synchronization edge occupy another priority queue. Hence, only one priority queue is left for $b_{1}$ and its synchronization edge. This argument repeats for all subsequent vertices of $S$.

### v3

@@ Appendix D Omitted Proofs of Section˜6
Refer to Figure˜6 for an illustration of our construction. We first show inductively that the edges on the left side are assigned to the same priority queues as their (adjacent) synchronization edges, and then we show the same for the edges on the right side.
Let $a_{1}$ be the edge with the leftmost left endpoint corresponding to a vertex $a\in S$. The k1k-1 heavy edges below $a_{1}$ require k1k-1 priority queues, none of which can be used by $a_{1}$ or its synchronization edge since both form a forbidden pattern with each of the heavy edges. Therefore, $a_{1}$ and its synchronization edge are assigned to the only remaining priority queue.
Let $b_{1}$ be the edge with the second leftmost left endpoint corresponding to a vertex $b\in S$. The k2k-2 heavy edges below $b_{1}$ require k2k-2 priority queues, none of which can be used by $b_{1}$ or its synchronization edge. Moreover, $a_{1}$ and its synchronization edge occupy another priority queue. Hence, only one priority queue is left for $b_{1}$ and its synchronization edge. This argument repeats for all subsequent vertices of $S$.